• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name hsmith

    Log entry time 11:28:32 on May 10, 2011

    Entry number 358789

    keyword=Left Q1 reset fault

    We have found a work around for the Left Q1 polarity switch. The polarity switch is not cycling properly due to one of possibly two problems, intermittent cam switch operation or radiation damaged to surface mount optocouplers on the control board that read the positive position of the switch. The switch has been moved manually to the negative position. Please do not try to change polarity because the motor has been disconnected and will generate a fault. The GUI displays polarity stuck on the control screen, this should have no effect on ramping the magnet to the desired current. Will investigate further at the end of the run.